OpenStudy (anonymous):

Thomas has $6.35 in dimes and quarters. The number of dimes is three more than three times the number of quarters. If q represents the number of quarters, then which of the following expressions represents the value of the number of dimes in cents?

OpenStudy (mayankdevnani):

x = quarters .....3x + 3 = dimes ....0.25x + 0.10(3x + 3) = 6.35 ....25x + 10(3x + 3) = 635 ....25x + 30x + 30 = 635 ....55x = 605 ....x = 11 quarters ....3x + 3 = 36 dimes